OpenStudy (anonymous):

The number of milligrams of a diabetic drug in a person's system after t hours is given by the function D(t)=20e^(-0.4t). Use the function to determine when the drug level will reach 0.01 mg (round to nearest hour).

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Now plug in D(t) = 0.01 and solve for t D(t)=20e^(-0.4t) 0.01 = 20e^(-0.4t) 0.01/20 = e^(-0.4t) 0.0005 = e^(-0.4t) ln(0.0005) = -0.4t -7.600902 = -0.4t -7.600902/(-0.4) = t 19.002255 = t t = 19.002255 Which rounds to 19. So it will take about 19 hours for the drug level to reach 0.01 mg

OpenStudy (anonymous):

I really appreciate you taking time to walk me through these, thank you so much! =)

jimthompson5910 (jim_thompson5910):

anytime